manufacturer wishes to determine the effectiveness of four types of machines
A, B, C and D in the production of bolts. For this, the number of defective bolts
produced by each machine on two alternative days are shown in the following table:
Day I Day II
Machine A 24 30
Machine B 41 44
Machine C 32 31
Machine D 28 38

Perform an analysis of variance method to determine whether there is a difference
between machines, at 5% level of significance.
1
Expert's answer
2020-06-01T17:07:25-0400

χ2="\sum"ri=1"\sum"cj=1=(Oij-Eij)2/Eij

Eij=RiCj/N, Ri - sum of i-th row, Cj - sum of j-th column, N - total sample size

E11=(54*125)/268=25, (O11-E11)2/E11=0.0416

E12=(54*143)/268=29, (O12-E12)2/E12=0.033

E21=(85*125)/268=40, (O21-E21)2/E21=0.024

E22=(85*143)/268=45, (O22-E22)2/E22=0.0227

E31=(63*125)/268=29, (O31-E31)2/E31=0.281

E32=(63*143)/268=34, (O32-E32)2/E32=0.29

E41=(66*125)/268=31, (O41-E41)2/E41=0.32

E42=(66*143)/268=35, (O42-E42)2/E42=0.2368

χ2=1.2491

degrees of freedom: (r-1)(c-1)=3

the critical value of chi square with 3df at 5% level significance: 7.82

χ2<critical value, so we accept that there is no difference between machines